### v1

# Shortcut Learning in In-Context Learning: A Survey

Abstract
Shortcut learning refers to the phenomenon where models employ simple, non-robust decision rules in practical tasks, which hinders their generalization and robustness. With the rapid development of large language models (LLMs) in recent years, an increasing number of studies have shown the impact of shortcut learning on LLMs. This paper provides a novel perspective to review relevant research on shortcut learning in In-Context Learning (ICL). It conducts a detailed exploration of the types of shortcuts in ICL tasks, their causes, available benchmarks, and strategies for mitigating shortcuts. Based on corresponding observations, it summarizes the unresolved issues in existing research and attempts to outline the future research landscape of shortcut learning.
Shortcut Learning in In-Context Learning: A Survey
Rui Song1,2, Yingji Li1,2, Fausto Giunchiglia1,3 Hao Xu1, 1College of Computer Science and Technology, Jilin University, Changchun, China, 2Key Laboratory of Symbol Computation and Knowledge Engineering of Ministry of Education, Jilin University, China, 3Department of Information Engineering and Computer Science, University of Trento, Italy Correspondence: xuhao@jlu.edu.cn
1 Introduction
In recent years, Large Language Models (LLMs) have emerged as a hotly pursued research direction, with the advent of major language models such as T5 Raffel et al. (2020), LLaMA Touvron et al. (2023), PaLM Chowdhery et al. (2023), GPT-3 Brown et al. (2020), Qwen2 Yang et al. (2024a), and GLM Zeng et al. (2023). LLMs demonstrate the ability of In-Context Learning (ICL), meaning they can learn from several demonstration examples within a given context Dong et al. (2022). However, an increasing number of studies have shown that LLMs are susceptible to shortcut learning in ICL and perform poorly in many aspects Du et al. (2024).
Shortcuts refer to decision rules that perform well on standard benchmarks but fail to transfer to more challenging testing conditions. Shortcut learning is the process by which a model exploits these decision rules during execution, leading to robust output results Geirhos et al. (2020). In the literature, shortcut learning is also referred to as superficial/spurious correlations Zhou et al. (2024c), the Clever Hans effect Lapuschkin et al. (2019), and various types of biases 111In our survey, bias refers to systematic errors within LLMs, rather than cultural or social biases that are studied in the field of LLM safety and fairness Zheng et al. (2024)., such as learning bias Du et al. (2024), label bias Fei et al. (2023), selection biases Wei et al. (2024), co-occurrence bias Kang and Choi (2023), and so forth. In practice, if LLMs establish shortcut ‘FlowersPosition’ based on given samples in Figure 1, predictions may fail when confronted with ‘Negative’ samples. Apart from this mapping from lexicon to specific labels, shortcuts also exist widely in various forms across different NLU tasks. For example, in Natural Language Inference (NLI) Zhou and Bansal (2020), Question-Answering (QA) Sen and Saffari (2020), and Reading Comprehension (RC) Lai et al. (2021), there is a preference for overlapping parts of two input branches. These unhealthy shortcuts have been widely proven to be harmful to the robustness, generalization, and fairness, and may lead to hallucinations in LLMs Li et al. (2023b); Ji et al. (2023).
Given the rapid development of LLMs and the impact of shortcut learning on them, our survey aims to systematically summarize, discuss, and consolidate the existing research, and to sensitize the community to the current advancements. Different from previous reviews on shortcut learning Ho et al. (2022); Du et al. (2024); Dogra et al. (2024); Bihani and Rayz (2024), our survey focuses on LLMs, while prior research primarily centered on Lightweight Pre-trained Language Models (LPLMs) such as BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019). Considering the differences in the training and application of LLMs, we deem it necessary to reassess the types of shortcuts, the causes of shortcut learning, benchmarks, and mitigation strategies, and to highlight the unresolved issues surrounding shortcut learning in LLMs. The organization of sections is as follows.
In Section 2, we summarize and generalize all the types of shortcuts that have been studied in LLMs. In Section 3, we summarize the possible causes of shortcuts in LLMs. In Section 4, we present the benchmarks commonly used in the evaluation tasks for shortcuts in LLMs. In Section 5, we provide mitigation strategies for different types of shortcuts. In Section 6, we conclude and discuss the existing research, and attempt to showcase potential directions for shortcut learning in LLMs. The detailed taxonomy is shown in Figure 2.
2 Types of Shortcuts
Compared to Du et al. (2024), we extend and summarize various types of shortcuts based on the characteristics of NLU tasks in LLMs. Specifically, we classify shortcuts into instinctive shortcuts and acquired shortcuts, based on whether LLMs rely on prompts to establish superficial correlations with labels or answers. But in reality, we believe that instinctive shortcuts can influence acquired shortcuts to some degree. In other words, instinctive shortcuts may facilitate LLMs in acquiring shortcuts from demonstrations Pan et al. (2023).
2.1 Instinctive Shortcuts
Instinctive shortcuts refer to the inherent preferences of LLMs for texts or patterns outside the examples provided in a given context. These preferences are not influenced by the contextual demonstrations but hinder the effective learning of the input-label relationships presented in those demonstrations. To avoid naming ambiguity, we have adopted the label bias definition proposed by Fei et al. (2023), which has been widely discussed.
Vanilla-label bias describes the model’s uncontextual preference for predicting certain label names. Zhao et al. (2021); Zheng et al. (2024) point out that regardless of the prompt, LLMs tend to assign higher probabilities to specific labels. Holtzman et al. (2021) demonstrates the existence of Surface Form Competition among different words, leading to a significant decline in prediction performance when low-frequency words are the correct answers. Similarly McCoy et al. (2023); Kang and Choi (2023) also confirm the preference of LLMs for high-frequency labels. Pan et al. (2023) show that replacing labels with meaningless words, such as ‘@#$’, leads to a decline in performance compared to using meaningful labels like ‘ABC’.
Context-label bias reveals the influence of context prompt, including factors such as sequence and format, on prediction outcomes. For example, subtle adjustments to the prompt format can lead to high instability in model performance, such as changing ‘Passage:<text>’ to ‘Passage <text>’ Sclar et al. (2024). Additionally, LLMs are also highly sensitive to term frequencies and task probabilities Razeghi et al. (2022). For instance, when asked to decode shift ciphers with various shift levels, GPT-4 scores 50% or above for the three most common shift numbers (1, 3, and 13), but scores below 3% for all other shift levels represented by different numbers McCoy et al. (2023). Moreover, LLMs are also highly sensitive to the order of inputs, whether it be the order of demonstrations or the order of candidate answer options. For example, LLMs are thought to favor options presented at a particular ranking position, such as first or last Wang et al. (2024); Pezeshkpour and Hruschka (2024); Zheng et al. (2024); Wei et al. (2024). While Lu et al. (2022) demonstrates that the sequence of presented samples can lead to a vast difference in performance, ranging from nearly achieving the current state-of-the-art to merely performing at the level of random guessing.
Domain-label bias describes the dependence on prior knowledge related to the task. Research has indicate that the in-domain words randomly sampled from the dataset are heavily biased towards the domain labels in ICL Fei et al. (2023). Accordingly, label words also often imply specific task types Si et al. (2023a). For example, the word ‘positive’ frequently appears in connection with sentiment classification tasks. If ‘positive’ were replaced with a meaningless symbol like ‘1’, LLMs would find it difficult to discern whether the task is sentiment classification or topic classification. This leads to LLMs relying on their pretrained demonstration semantic priors in ICL predictions, rather than learning from the input-label relationships presented in demonstrations Jang et al. (2024). This prior can also be observed in QA tasks, where LLMs are able to make correct predictions even without being given options Balepur et al. (2024).
Apart from the semantic priors between texts and labels, entity knowledge within LLMs can also lead to bias Wang et al. (2022). For instance, LLMs may assume that ‘Bill Gates’ is the founder of ‘Microsoft’, even when given the context ‘Bill Gates went to Microsoft Building 99’ where the relationship between ‘Bill Gates’ and ‘Microsoft’ is visitor. Similarly, in relation extraction tasks where the intended relationship to be extracted is unrelated to the existing relationships in the input, LLMs may still extract these unrelated relationships Zhang et al. (2024b). In multi-hop QA tasks, LLMs tend to directly provide answers by leveraging potential associations between the head entity and the entity to be predicted, rather than engaging in correct reasoning paths Li et al. (2024a); Ju et al. (2024).
Reason bias is a newly summarized shortcut type in this paper compared with Fei et al. (2023). It indicates that LLMs tend to get answers directly from the provided examples rather than reasoning the underlying patterns Ali et al. (2024), especially in complex reasoning and summarization tasks involving long contexts Li et al. (2024b); Wu et al. (2024); Yang et al. (2024c). At first glance, reason bias may seem to rely on demonstrations, but essentially, it represents a tendency of LLMs to take shortcuts when solving complex problems. Therefore, we categorize it as an instinctive shortcut.
2.2 Acquired Shortcuts
Acquired shortcuts refer to the shortcuts present in demonstrations, which LLMs capture as non-robust patterns through observation of these demonstrations. These patterns are then utilized by LLMs for inference, leading to a decline in performance.
Lexicon. When spurious correlation are established between certain lexical features and specific labels, the corresponding lexicon may then become a shortcut. Arbitrary text can serve as lexicon shortcuts, including letters, signs, numbers, negation words, adverbs of degree, or even recurring sentences Tang et al. (2023). Among them, negation words have garnered the most attention from researchers, as LLMs often provide contradictory predictions due to these words Sun et al. (2024); Yuan et al. (2024); Zhou et al. (2024d) But studies have indicated that LLMs are more likely to rely on n-grams and content words rather than others, such as stop words Si et al. (2023b); Pacchiardi et al. (2024).
Concept. Concept shortcut refers to the phenomenon where the concepts of certain texts in the demonstrations are disproportionately associated with the labels Zhou et al. (2024c). For example, if sentences containing cities are associated with negative emotions and samples containing countries are also associated with negative emotions, when LLMs encounter new city words in the test sample, they may predict it as having negative emotions Zhou et al. (2024d).
Overlap. Overlap shortcuts often emerge in tasks that involve two branches, such as NLI and QA. In NLI, overlap between the premise and the hypothesis can be a strong indicator of Entailment Sun et al. (2024); Yuan et al. (2024). In extractive QA, words that appear in both the question and the background may be used as anchors. LLMs tend to extract answers around these anchors to reduce the error rate Levy et al. (2023).
Position. LLMs use positional information as a basis for prediction. For example, if the answer is located at the beginning of the text, the model is less likely to make an error Levy et al. (2023). LLMs can also briefly read the opening segment to predict the answer without needing to read the entire document Li et al. (2024b). Conversely, Yuan et al. (2024) show that adding content unrelated to the answer at the beginning of a sentence can degrade the prediction performance of LLMs. Additionally, Zhao et al. (2021) also observe that answers closer to the end of the prompt are more likely to be repeated by the model.
Text Style. The writing style of the text is also a possible shortcut Qi et al. (2021). Tang et al. (2023); Yuan et al. (2024) demonstrate that shifts in presentation style can have an impact on performance, such as when standard English is rewritten into Shakespearean English. LLMs can also capture the correlation between style and labels, leading to a decline in the generalization ability of ICL Zhou et al. (2024d).
Group Dynamics. It refers to the interactions, relationships, and influences among members within a group Cartwright and Zander (1968). Similarly, in the decision-making process, LLMs may evaluate a choice based on other available options. For instance, if other options are even numbers, the model may tend to lean towards an odd number choice Balepur et al. (2024). This type of shortcut can also be associated with the A-Not-B error in the field of cognitive science Popick et al. (2011). If ‘A’ appears more frequently in the demonstrations, the model is more inclined to predict ‘A’, even if ‘B’ is correct Gupta et al. (2023); Han et al. (2024).
3 Causes of Shortcut Learning
In light of the impact of shortcut learning, some methods attempt to provide explanations for the causes of shortcut learning in ICL. This paper divides the causes of shortcut learning in LLMs into three parts: LLMs training, skewed demonstrations, and LLMs size.
3.1 LLMs training
Instinctive shortcuts reveal the inherent biases of LLMs. The causes of inherent biases are diverse and closely related to the training process of LLMs, involving factors such as data distribution and instruction tuning. This makes it difficult for LLMs to fully overcome the prediction biases acquired from pre-training data and to consider all contextual information equally during the ICL process Kossen et al. (2024).
Pretraining. Some studies suggest that instinctive shortcuts originate from the data distribution during the pretraining process Razeghi et al. (2022); McCoy et al. (2023), which leads to a bias in the predictions of LLMs towards high-frequency and frequently co-occurred words Kang and Choi (2023); Ju et al. (2024). Rare behavioral patterns in the training data, such as converse relations, can also confuse LLMs, causing them to rely on shortcuts when making judgments Qi et al. (2023). Furthermore, Zhou et al. (2024d) utilize SHAP analysis Lundberg and Lee (2017) to confirm that high-frequency words that have spurious correlations with a certain label contribute more significantly to the model’s predictions for that label. Additionally, if the frequency of a particular LLMs task is low, its prediction results will also be inferior compared to those of high-frequency tasks McCoy et al. (2023). Based on the observations mentioned above, Han et al. (2022); Fei et al. (2023) implicitly assume that there is a shift in the label marginal of the ICL predictive distribution, and Jiang et al. (2023) theoretically proves that the emergence of shortcuts stems from label shifting caused by the contextual label marginal typically deviating from the true label distribution.
Instruction Tuning. Another perspective holds that the instinctive shortcuts in LLMs originate from the process of instruction tuning, during which the LLMs learn spurious correlations within the instructions Sun et al. (2024). In complex multitasking scenarios, LLMs can also learn more intricate associations between tasks, features, and labels during the fine-tuning process. As a result, LLMs can exploit these spurious correlations simply by being provided with useful task information Pan et al. (2023).
3.2 Skewed demonstrations
Corresponding to LLMs training, skewed demonstrations are primarily responsible for the generation of acquired shortcuts, as LLMs learn from and reason based on the models presented in the demonstrations. If the demonstrations are skewed, the corresponding reasoning results will also be affected. Many studies have confirmed this by adding specific shortcut triggers to the demonstrations, where the shortcuts in the demonstrations caused a decline in the performance of LLMs Tang et al. (2023); Si et al. (2023b); Levy et al. (2023).
3.3 LLMs size
While larger LLMs exhibit exceptional capabilities in semantic understanding and reasoning, numerous studies have shown that increasing the LLMs’ scale does not mitigate shortcut learning Kang and Choi (2023); Zhou et al. (2024d); Sclar et al. (2024). Rather, even larger LLMs are more prone to exploiting shortcuts, whether they are instinctive shortcuts or acquired shortcuts Razeghi et al. (2022); Tang et al. (2023); Balepur et al. (2024); Yuan et al. (2024); Han et al. (2024). Although not directly stated, Pan et al. (2023) provide some explanation for this phenomenon: larger LLMs can utilize more demonstrations to enhance their ability to learn new input-label mappings, whereas smaller LLMs cannot. Consequently, when there are more shortcuts in the context, larger LLMs are more susceptible to being influenced by these shortcuts.
4 Benchmarks of Shortcut Learning
ICL can encompass many of NLU tasks, targeting different types of shortcuts. Therefore, we deem it necessary to show existing evaluation benchmarks to support related research. Specifically, we summarize the datasets used in existing research, along with the corresponding quantitative metrics.
4.1 Datasets
Most studies have not developed shortcut evaluation datasets, but instead utilize existing datasets Zhao et al. (2021); Fei et al. (2023); Gupta et al. (2023) or make simple modifications to adapt them to their own shortcut evaluation tasks. The most common approach is to add or delete elements from the samples in existing datasets to construct samples that contain shortcuts. For exampple, Tang et al. (2023); Si et al. (2023b) introduce shortcuts by injecting shortcut trigger words, typically a sequence of characters, into the samples, and used this method to evaluate the impact of these shortcuts on ICL. Alternatively, shuffling or deleting option IDs can be used to evaluate the impact of option order on ICL Zheng et al. (2024). Some methods also select a subset of samples that meet certain criteria from existing datasets for experimentation. For example, Sun et al. (2024) choose samples containing words like ‘film’ or ‘movie’ for lexicon shortcuts. LLMs can also play a role for dataset modifications, such as extracting concepts from text for concept shortcuts Zhou et al. (2024c).
There are also methods that propose corresponding benchmarks for different types of shortcuts to better evaluate them, as the the datasets mentioned above are not specifically designed for shortcut evaluation. Shortcut Maze is a comprehensive benchmark for text classification that categorizes shortcuts into lexicon, style, and concept Zhou et al. (2024d). Similarly, Shortcut Suite includes lexicon, overlap, position, and style shortcuts, but it focuses its research on NLI tasks Yuan et al. (2024). Apart from classification tasks, QA benchmarks are also a common scenario for exploring shortcuts. ShortcutQA performs counterfactual editing of samples by adding or excluding shortcut triggers based on the confidence of the target model to obtain shortcut-related samples. It involves various types of shortcuts, including entity, overlap, position, and more Levy et al. (2023). EUREQA is a complex multi-hop QA dataset designed to reduce semantic relevance by adding deceptive semantic cues that distract attention and are unrelated to the correct answer, and to measure the ability of LLMs to perform extensive chained reasoning processes Li et al. (2024a). MMLU-Pro+ introduces various interfering shortcuts into the correct options of reasoning questions across 14 different domains to enhance its ability to evaluate high-order reasoning skills in LLMs Taghanaki et al. (2024). ReWild focuses more on testing LLMs at a fine-grained level in various aspects such as mathematics, logic, and common sense, including the selection and execution of strategies, as well as the tendency to adopt undesirable shortcuts Yang et al. (2024c). ConvRe systematically evaluates the ability of LLM to recognize and process inverse relations, determining whether these inverse relations are affected by shortcuts through the mutual matching of knowledge triples with their corresponding textual descriptions Qi et al. (2023).
4.2 Metrics
The most commonly used metrics in shortcut evaluation are general evaluation methods for NLP tasks such as accuracy, AUC score Chen et al. (2023) as well as F1 score in classification tasks Tang et al. (2023); Fei et al. (2023), Hits@n in retrieval related tasks Kang and Choi (2023); Pezeshkpour and Hruschka (2024), and exact match score in machine reading comprehension tasks Zhang et al. (2024a); Wang et al. (2023a). If the degree of shortcut learning in ICL is significant, then the corresponding perfomance will also undergo changes Zhou et al. (2024c); Han et al. (2024); Zhou et al. (2024d). Similarly, Fluctuation Rate is designed to represent the ratio of the change in prediction results to the total after a sample is perturbed Wei et al. (2024). In other studies, Conflict Rate is also equivalent to Fluctuation Rate Zheng et al. (2024). Taghanaki et al. (2024) define the Shortcut Selection Ratio, which is used to quantify the tendency of LLMs to maintain their original choices after introducing distracting items.
5 Mitigation of Shortcut Learning
Given the widespread existence and dangers of shortcuts, numerous studies have investigated corresponding methods to mitigate them. Based on the differences in the targets addressed by these mitigation methods, we categorize the existing approaches into three types: data-centric, model-centric, and prompt-centric.
5.1 Data-centric Approach
As summarized in Section 3, the shortcuts in LLMs are largely related to the distribution of training data. Therefore, an intuitive approach is to sample task-relevant examples to obtain balanced training data. Kang and Choi (2023) alleviate the dependence on shortcuts by filtering out samples with high co-occurrence probabilities and guiding the retraining of LLMs. Zhou et al. (2024c) employ counterfactual samples generated by LLMs for upsampling and re-trains to mitigate the reliance on shortcuts. Nakada et al. (2024) proposes a systematic oversampling method that leverages the capabilities of LLMs to generate high-quality synthetic data for minority groups. However, the cost of fine-tuning LLMs is too high, and it can also result in catastrophic forgetting Li and Lee (2024). Therefore, there are relatively few corresponding studies.
5.2 Model-centric Approach
The model-centric approach mitigates the impact of shortcuts on the results by manipulating neurons and the predictive probability distribution. The main methods can be divided into those based on model pruning and those based on calibration.
Model pruning alleviates the impact of shortcuts on the model itself by removing biased neurons. This operation can be supported by research on the relationship between neurons and knowledge storage, where neurons are the basic units for storing knowledge and patterns in LLMs Zhao et al. (2024a). In this process, some commonly used post-hoc interpretable methods, such as Integrated Gradients Sundararajan et al. (2017), are used to locate the neurons activated during shortcut learning. By removing these neurons, LLMs can be prompted to explore alternative and correct reasoning paths Ju et al. (2024); Yang et al. (2024b). Generally, the weights of neurons with high shortcut scores are set to 0 to achieve the purpose of pruning Ali et al. (2024). By quantifying the contribution of feedforward neural networks and attention heads to prediction results, Zhou et al. (2024b) identify biased components of LLMs and alleviates their prompt brittleness by removing the corresponding components. However, the negative impact of such component removal on LLMs is still debatable, as neurons have been observed to be polysemous, meaning that a single neuron can be activated by multiple unrelated terms Liu et al. (2022); Bills et al. (2023). That means when shortcut components are removed, some corresponding useful knowledge may also be discarded, introducing new noise.
Calibration corrects biased outputs of LLMs by modifying theirs probability distribution Guo et al. (2017). Contextual Calibration obtains context-driven biased distributions by replacing the sample to be predicted with meaningless inputs, such as ‘N/A’ Zhao et al. (2021). In an ideal unbiased environment, even if the sample to be predicted is set to ‘N/A’, LLMs (Large Language Models) should yield an output similar to 50% Positive and 50% Negative. If the prediction distribution shifts, then this shift is attributed to the corresponding context. Subsequently, during actual predictions, LLMs’ outputs are adjusted based on the biased distribution to mitigate biases in the context.
As research progresses gradually, Contextual Calibration has spawned various calibration methods tailored for different shortcut mitigation scenarios. Prototypical Calibration employs a Gaussian mixture distribution to estimate prototypical clusters for all categories and then utilizes the likelihood of these prototype clusters to calibrate LLMs’ predictions Han et al. (2022). Domain-context Calibration employs randomly selected in-domain words from the task corpus to estimate the label bias of the language model and utilizes this estimation for bias mitigation Fei et al. (2023). Similarly, PMIDC re-weighs scores based on the likelihood of hypotheses (answers) given premises (questions) in a specific task domain Holtzman et al. (2021).
But these calibration schemes may fail due to their inability to effectively estimate contextual bias with only content-free and in-domain random tokens. To address this, Batch Calibration updates the bias as more batches of input data are processed, allowing the bias to stabilize after multiple mini-batches are considered Zhou et al. (2024a). Generative Calibration adjusts the margin of the labels estimated via Monte-Carlo sampling over the in-context model to simply calibrate the contextual prediction distribution Jiang et al. (2023). In-Context Calibration calibrates the test distribution of the demonstration using the prior semantics expected from the samples Jang et al. (2024). PriDe is applied to the task of multiple choice questions by arranging the option content on a small number of test samples to estimate prior information, and then using the estimated prior information to debias the remaining samples Zheng et al. (2024). Unlike the aforementioned probability-based calibration methods, NOISYICL calibrates language models by adding Gaussian noise to the parameters of LLMs, thereby reducing the prediction bias and unreliable confidence in ICL Zhao et al. (2024b).
5.3 Prompt-centric Approach
The prompt-centric approach aims to reduce LLMs’ focus on shortcuts by modifying the prompts. Depending on the specific part of the prompt that is modified, it can be classified into three methods: shortcut-based method, instruction format-based method, and prompt search-based method.
Shortcut-based method mitigates the unhealthy focus of LLMs on shortcuts by masking or perturbing shortcut words or concepts. Zhou et al. (2024c) masks words with high relevance to shortcut concepts, where relevance is measured using Pointwise Mutual Information. Wang et al. (2023a) replace entities that may rely on shortcuts with placeholders and alleviates the focus on a particular entity by guiding LLMs to understand the placeholders as a set of similar entities.
Instruction format-based method guides LLMs to produce unbiased responses by modifying the format of the prompt or adding more detailed instructions. For example, it is advisable to utilize semantic verbalizers as much as possible to ensure that the label tokens have semantic relevance to the task Si et al. (2023a). Alternatively, in multiple-choice scenarios, the multiple options can be rearranged, and the rearranged results can be integrated using a majority voting method from LLMs to mitigate the risks associated with shortcuts related to option order Pezeshkpour and Hruschka (2024); Wang et al. (2024). Sun et al. (2024) recommend avoiding overloading the LLMs with too much task information that could lead to remembering previous false associations by mixing prompts and labels. Mixing prompts is achieved by substituting the first sample in the demonstration for mixed prompts, while mixed labels involves randomly replacing the labels with meaningless symbolic tags. In addition to the aforementioned modifications to prompt formats, some methods also use hints Qi et al. (2023) or Chain-of-Thought (CoT) Yuan et al. (2024) to guide LLMs in making more reasonable inferences. Zhang et al. (2024a) utilize the chain of thought generated by LLMs as a mediator variable, and adjust the causal relationship between the input prompt and the output answer through front-door adjustment to mitigate ICL bias.
Prompt search-based method aims to mitigate biases acquired by LLMs from context by searching for unbiased prompts. Unlike instruction format-based method, this approach focuses on retrieving reasonable prompts within a certain range, rather than having a clear modification target. Prompt search-based method can easily be associated with some recent retrieval-augmented approaches, which enhance the quality of ICL by retrieving samples from the corpus that are helpful for predictions Yoran et al. (2024); Ram et al. (2023). Chen et al. (2023) observe a strong negative correlation between the sensitivity and accuracy of ICL, indicating that predictions sensitive to perturbations are unlikely to be correct. Based on this, SENSEL is proposed to avoid providing users with incorrect predictions by discarding samples that are sensitive to perturbations. Lu et al. (2022) utilize the predicted label distribution statistics and propose an entropy-based metric to measure the quality of candidate prompts with diffenret order. Gonen et al. (2023) assumes that lower perplexity correlates with better performance and selects prompts with low perplexity as input.
6 Rethinking of Shortcut learning
This section compares the shortcut learning in ICL with that in LPLMs, and outlines several potential future research directions.
6.1 Differences from Shortcuts in LPLMs
Given that previous reviews have primarily focused on research into LPLMs, we deem it necessary to elucidate the similarities and differences between shortcut learning in LLMs and LPLMs.
The difference between them primarily arises from the distinctions in research paradigms between LPLMs and LLMs. LPLMs typically require fine-tuning for different tasks, whereas LLMs unify various NLU tasks and only need a few samples to guide their inference. Therefore, ICL encounters a broader range of shortcut types compared to those in LPLMs, even though both include shortcut types such as lexicon Niven and Kao (2019), overlap Lai et al. (2021), position Ko et al. (2020), and text style Qi et al. (2021). For instance, since LLMs do not require fine-tuning, the biases inherent in LLMs themselves can strongly influence various tasks, leading to the emergence of the instinctive shortcuts. Furthermore, since LLMs can easily infer answers from few samples, they naturally encounter shortcuts arising from these samples’ group dynamics. The shortcuts related to group dynamics may, to some extent, resemble the challenges faced in long-tailed classification, where models tend to learn answers from head classes rather than tail classes Li et al. (2023a). But besides the skew in categories, LLMs are also influenced by more complex group decision-making process, such as the parity of options Balepur et al. (2024). This can be attributed to LLMs’ stronger comprehension abilities, enabling them to uncover complex correlations that LPLMs are unable to detect.
Another notable difference lies in shortcut mitigation strategies. For LPLMs, a widely adopted mitigation approach involves adjusting the model during fine-tuning, which encompasses techniques such as adversarial training Rashid et al. (2021), regularization Stacey et al. (2022), Product-of-Experts Sanh et al. (2021), reweighting Liu et al. (2021), contrastive learning Song et al. (2023), and other related operations. For LLMs, fine-tuning to mitigate shortcuts is rare, and the primary methods focus on research based on forward propagation. But we can still find extensions of LPLMs in the related research on LLMs. For instance, Calibration methods are similar to Product-of-Experts, both of which adjust the prediction probabilities.
6.2 Future Directions
More Robust Evaluation Benchmark. Siska et al. (2024) demonstrate that the correlation between the performance exhibited by the model and the prompts is significant, which suggests that the evaluation benchmark itself also exhibits bias. When the model responds to certain similar prompts within the benchmark, it tends to produce similar results. Pacchiardi et al. (2024) also indicate that LLMs can exploit shortcuts in the benchmark to obtain inflated evaluation scores. In addition, data contamination and leakage can also impact the evaluation of LLMs’ shortcuts, as LLMs tend to provide good prediction results directly for benchmarks that have appeared in the training corpus Zhou et al. (2023). Balepur et al. (2024) also imply that leakage of the test set might have an impact on multiple-choice questions without the question. The aforementioned studies suggest that caution should be exercised when interpreting the performance results of LLMs, and the impact of shortcuts should be fully considered. Furthermore, efforts should be made to minimize the potential influence of shortcuts on prediction results before constructing benchmarks.
More Shortcut Related Tasks. Despite the extensive exploration of shortcut learning in ICL, given the diversity of NLP tasks, it remains crucial to expand the boundaries of research tasks. For instance, in Table QA tasks, only a small portion of the entire table is relevant to deriving the answer to a given question, and irrelevant information may become potential shortcuts that affect the prediction outcomes of LLMs Patnaik et al. (2024). These expanded tasks are highly dependent on relevant benchmarks, so designing and constructing corresponding datasets and evaluation methods for new tasks also require attention.
More Interpretability. Although the studies in Section 3 explore the causes of shortcuts from different aspects, most of them are empirical summaries based on experimental observations, lacking an interpretation of the causes of shortcuts from an explainable perspective. For instance, how ICL influences the role of instinctive shortcuts in different NLU tasks, and at which step of the forward propagation in LLMs do these shortcuts exert their impact, is it during aggregation or distribution Wang et al. (2023b)? Can self-explanation methods of LLMs be effective in this process Huang et al. (2023)? Interpretable methods can enhance the intuitive presentation of shortcut learning in ICL and facilitate effective mitigation of shortcuts by addressing their underlying causes.
More Discussion in Shortcut Unknown Scenarios. In the research of LPLMs, shortcut identification is a crucial preliminary step for shortcut mitigation Du et al. (2024), as in many cases, the types of shortcuts in tasks are unknown, or multiple shortcuts coexist. But in ICL, most studies assume that LLMs are vulnerable to shortcuts, or simply presume that the types of shortcuts are known and explore methods to assess performance under these known shortcuts. This idealized and strong assumption of known shortcuts may fail when facing new tasks, leading to limitations in shortcut mitigation solutions. Therefore, exploring how LLMs identify and utilize shortcuts in tasks where the shortcuts are unknown is also one of the subsequent research directions.
Decoupling of Instinctive and Acquired Shortcuts. Based on the discussion in Section 3, instinctive shortcuts are established during the training process of LLMs, while acquired shortcuts originate from demonstrations in the inference process. Instinctive shortcuts should emerge earlier in the lifecycle of LLMs. Consequently, research solely focused on acquired shortcuts may be potentially influenced by instinctive shortcuts, leading to biased conclusions, as discussed by Li et al. (2023b) in LLMs bais. Therefore, there is a need to strengthen the exploration of how these two different types of shortcuts interact with each other. Furthermore, in the research on analyzing and mitigating acquired shortcuts, it is important to consider reducing or eliminating the influence of instinctive shortcuts to ensure the objectivity of the experimental results.
7 Conclusion
This paper comprehensively reviews the existing literature related to shortcut learning in LLMs, conducts a thorough classification of shortcuts in ICL, discusses their causes, presents existing evaluation benchmarks and mitigation strategies, and identifies key challenges and potential directions for future work. To our knowledge, this is the first review dedicated to the phenomenon of shortcut learning in ICL. We aim to emphasize the current research status of shortcut learning in the context of LLMs and provide insights to guide future work.